OpenStudy (anonymous):

a patient is to take 4 2/3 tablespoons of medicine in 2 equally daily doses. how much medicine is to be taken in each dose

OpenStudy (anonymous):

4 2/3 = 14/3 (14/3) / 2 = 14/6 = 2 2/6 = 2 1/3 or you may : (4 2/3) / 2 = 2 1/3